F(x)= x^2 , g(x)= 2x+1
(gof)(-3)
When I did this problem I got a different answer than what it says . I need help on what I I possibly did wrong.
(Fog)(1/2)

Answer by Boreal(15235)

g (f (x))=2x^2+1
The f(x) goes where the x is in the g(x) function
(g o f)(-3)=19
f(-3)=9
g(9)=19
(f o g) is (2x+1)^2 or for (1/2) would be 4
4x^2+4x+1
1+2+1=4
g(1/2)=2
f(2)=4